Grade 12th passAlgebra

Write a Pythagorean triplet whose one member is.
6.......

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

For any natural number m, we know that 2m, m2–1, m2+1 is a Pythagorean triplet.
2m=6
⇒m= 6/2
=3
m2–1=
32–1
=9–1
=8
m2+1= 32+ 1
= 9+1
= 10
∴(6, 8, 10) is a Pythagorean triplet.

Thanks